Direct Labor Variances

Greeson Clothes Company produced 24,000 units during June of the current year. The Cutting Department used 4,600 direct labor hours at an actual rate of $14.70 per hour. The Sewing Department used 7,600 direct labor hours at an actual rate of $14.40 per hour. Assume that there were no work in processinventories in either department at the beginning or end of the month. The standard labor rate is $14.60. The standard labor time for the Cutting and Sewing departments is 0.20 hour and 0.30 hour per unit, respectively.

a. Determine the direct labor rate, direct labor time, and total direct labor cost variance for the Cutting Department and Sewing Department. Enter a favorable variance as a negative number using a minus sign and an unfavorable variance as a positive number.

Cutting Department Sewing Department
Direct labor rate variance $   $  
Direct labor time variance $   $  
Total direct labor cost variance $   $  
Answer #1
Cutting Department:
Direct labor rate variance 460 Unfavorable =4600*(14.7-14.6)
Direct labor time variance -2920 Favorable =14.6*(4600-24000*0.2)
Total direct labor cost variance -2460 Favorable =(4600*14.7)-(24000*0.2*14.6)
Sewing Department:
Direct labor rate variance -1520 Favorable =7600*(14.4-14.6)
Direct labor time variance 5840 Unfavorable =14.6*(7600-24000*0.3)
Total direct labor cost variance 4320 Unfavorable =(7600*14.4)-(24000*0.3*14.6)